HIGH RESOLUTION INFRARED SPECTROSCOPY OF THE PO_{2} RADICAL

\underline{MICHAEL A. LAWSON}, KRISTIAN J. HOFFMAN and PAUL B. DAVIES, Department of Chemistry, University of Cambridge, Lensfield Road, Cambridge, CB2 1EW, U.K..

The infrared absorption spectrum associated with the asymmetric stretching fundamental of PO_{2} (v_{3}) has been measured using high resolution tunable diode laser spectroscopy. The free radical was formed in large concentrations in a fast-flow system by reacting white phosphorus vapour with atomic oxygen. More than 700 new lines have been assigned between 1290 and 1350 cm^{−1}. Data was combined with infrared and LMR lines, obtained from previous studies and simultaneously fitted using the CALPGM suite. The precise constants derived definitively exclude a series of unassigned transitions from belonging to the v_{3} fundamental band and potential assignments for that series are discussed.